OpenStudy (kidrah69):

For math if i wanted to type 4timesX how would i type that in LaTex to make it less complicated since x <---multiplacation X<------x looks the same

OpenStudy (nobrainiachere):

Use an asterisk? o.o

OpenStudy (nobrainiachere):

Like this, \(\Huge{4{*}X}\)

OpenStudy (kidrah69):

I forgot about that xD

OpenStudy (kidrah69):

ok anyways thanks :)

OpenStudy (nobrainiachere):

:/

OpenStudy (kidrah69):

LAG >:(

OpenStudy (nobrainiachere):

np

OpenStudy (unklerhaukus):

`\[4\times x\]` \[4\times x\]

TheSmartOne (thesmartone):

Or you can do \(\sf 4x\) Or \(\sf 4(x)\) :P

OpenStudy (unklerhaukus):

or `\[4\cdot x\]` \[4\cdot x\]